### v1

@@ 3.3.1 Latent video clip patch embedding
Uniform frame patch embedding. As illustrated in Fig. 3 (a), we apply the patch embedding technique outlined in ViT [12] to each video frame individually. Specifically, $n_{f}$, $n_{h}$, and $n_{w}$ are equivalent to $F$, $\frac{H}{h}$, and $\frac{W}{w}$ when non-overlapping image patches are extracted from every video frame. Here, $h$ and $w$ denote the height and weight of the image patch, respectively.
@@ 3.3.2 Timestep-class information injection
From simple and direct integration to complex and nuanced integration perspective, we explore two methods for integrating timestep or class information $c$ into our model. The first approach involves treating it as tokens, and we refer to this approach as all tokens. The second method is akin to adaptive layer normalization (AdaLN) [40, 39]. We employ linear regression to compute $\gamma_{c}$ and $\beta_{c}$ based on the input $c$, resulting in the equation $AdaLN(h,c)=\gamma_{c}\text{LayerNorm}(h)+\beta_{c}$, where $h$ represents the hidden embeddings within the Transformer blocks. Furthermore, we also perform regression on $\alpha_{c}$, which is applied directly before any residual connections (RCs) within the Transformer block, resulting in $RCs(h,c)=\alpha_{c}h+AdaLN(h,c)$. We refer to this as scalable adaptive layer normalization (S-AdaLN). The architecture of S-AdaLN is shown in Fig. 4a.
@@ 3.3.3 Temporal positional embedding
Temporal positional embedding enables a model to comprehend the temporal signal. We explore two methods as follows for injecting temporal positional embedding into the model: 1) the absolute positional encoding method incorporates sine and cosine functions with varying frequencies [64] to enable the model to recognize the precise position of each frame within the video sequence; 2) the relative positional encoding method employs rotary positional embedding (RoPE) [60] to enable the model to grasp the temporal relationships between successive frames.
@@ 3.3.4 Enhancing video generation with learning strategies
Learning with pre-trained models. The pre-trained image generation models have learned what the world looks like. Thus, there are many video generation works that ground their models on pre-trained image generation models to learn how the world moves [69, 3]. However, these works mainly build on U-Net within latent diffusion models. The necessity of Transformer-based latent diffusion models is worth exploring.
We initialize Latte from a pre-trained DiT model on ImageNet [39, 9]. Directly initializing from the pre-trained DiT model will encounter the problem of missing or incompatible parameters. To address these, we implement the following strategies. In pre-trained DiT, a positional embedding $\boldsymbol{p}\in\mathbb{R}^{n_{h}\times n_{w}\times d}$ is applied to each token. However, in our video generation model, we have a token count that is $n_{f}$ times greater than that of the pre-trained DiT model. We thus temporally replicate the positional embedding $n_{f}$ times from $\boldsymbol{p}\in\mathbb{R}^{n_{h}\times n_{w}\times d}$ to $\boldsymbol{p}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}\times d}$. Furthermore, the pre-trained DiT includes a label embedding layer, and the number of categories is 1000. Nevertheless, the used video dataset either lacks label information or encompasses a significantly smaller number of categories in comparison to ImageNet. Since we target both unconditional and class-conditional video generation, the original label embedding layer in DiT is inappropriate for our tasks, we opt to directly discard the label embedding in DiT and apply zero-initialization.
Learning with image-video joint training. The prior work on the CNN-based video diffusion model proposes a joint image-video training strategy that greatly improves the quality of the generated videos [18]. We explore whether this training strategy can also improve the performance of the Transformer-based video diffusion model. To implement simultaneous training for video and image generation, We append randomly selected video frames from the same dataset to the end of the chosen videos and each frame is independently sampled. In order to ensure our model can generate continuous videos, tokens related to video content are used in the temporal module for modeling temporal information, while frame tokens are excluded.
@@ 4 Experiments
Table 1: Inception Score and FID comparisons of Latte against other state-of-the-art on the UCF101 and FaceForensics datasets, respectively. We use the pre-trained models provided by PVDM to generate corresponding videos and report their correspondence values. Here, “IMG” means video-image joint training.
Method FaceForensics SkyTimelapse UCF101 Taichi-HD
MoCoGAN 124.7 206.6 2886.9 -
VideoGPT 185.9 222.7 2880.6 -
MoCoGAN-HD 111.8 164.1 1729.6 128.1
DIGAN 62.5 83.11 1630.2 156.7
StyleGAN-V 47.41 79.52 1431.0 -
PVDM 355.92 75.48 1141.9 540.2
MoStGAN-V 39.70 65.30 1380.3 -
LVDM - 95.20 372.0 99.0
Latte (ours) 34.00 59.82 477.97 159.60
Latte+IMG (ours) 27.08 42.67 333.61 97.09
Table 2: FVD values of video generation models on different datasets. FVD values for other baseline models are reported and sourced from the reference StyleGAN-V or the original paper. Additionally, we use the official code of PVDM, strictly adhere to the training method, retrain on FaceForensics and TaichiHD, and report their FVD results. Meanwhile, we use the pre-trained models provided by PVDM on UCF101 and SkyTimelapse to generate corresponding videos and report their FVD values. Here, “IMG” means video-image joint training.

### v2

@@ 3.3.1 Latent video clip patch embedding
Uniform frame patch embedding. As illustrated in Fig. 3 (a), we apply the patch embedding technique outlined in ViT Dosovitskiy et al. (2021) to each video frame individually. Specifically, $n_{f}$, $n_{h}$, and $n_{w}$ are equivalent to $F$, $\frac{H}{h}$, and $\frac{W}{w}$ when non-overlapping image patches are extracted from every video frame. Here, $h$ and $w$ denote the height and width of the image patch, respectively.
@@ 3.3.2 Timestep-class information injection
From simple and direct integration to complex and nuanced integration perspective, we explore two methods for integrating timestep or class information $c$ into our model. The first approach involves treating it as tokens, and we refer to this approach as all tokens. The second method is akin to adaptive layer normalization (AdaLN) following DiT Perez et al. (2018); Peebles & Xie (2023). We employ linear regression to compute $\gamma_{c}$ and $\beta_{c}$ based on the input $c$, resulting in the equation $AdaLN(h,c)=\gamma_{c}\text{LayerNorm}(h)+\beta_{c}$, where $h$ represents the hidden embeddings within the Transformer blocks. Furthermore, we also perform regression on $\alpha_{c}$, which is applied directly before any residual connections (RCs) within the Transformer block, resulting in $RCs(h,c)=\alpha_{c}h+AdaLN(h,c)$. We refer to this as scalable adaptive layer normalization (S-AdaLN). The architecture of S-AdaLN is shown in Fig. 10 of the Appendix.
@@ 3.3.3 Temporal positional embedding
Temporal positional embedding enables a model to comprehend the temporal signal. We explore two methods as follows for injecting temporal positional embedding into the model: 1) the absolute positional encoding method incorporates sine and cosine functions with varying frequencies Vaswani et al. (2017) to enable the model to recognize the precise position of each frame within the video sequence; 2) the relative positional encoding method employs rotary positional embedding (RoPE) Su et al. (2024) to enable the model to grasp the temporal relationships between successive frames.
@@ 3.3.4 Enhancing video generation with learning strategies
Learning with pre-trained models. The pre-trained image generation models have learned what the world looks like. Thus, there are many video generation works that ground their models on pre-trained image generation models to learn how the world moves Wang et al. (2024a); Blattmann et al. (2023a). However, these works mainly build on U-Net within latent diffusion models. The necessity of Transformer-based latent diffusion models is worth exploring.
We initialize Latte from a pre-trained DiT model on ImageNet Peebles & Xie (2023); Deng et al. (2009). Directly initializing from the pre-trained DiT model will encounter the problem of missing or incompatible parameters. To address these, we implement the following strategies. In pre-trained DiT, a positional embedding $\bm{p}\in\mathbb{R}^{n_{h}\times n_{w}\times d}$ is applied to each token. However, in our video generation model, we have a token count that is $n_{f}$ times greater than that of the pre-trained DiT model. We thus temporally replicate the positional embedding $n_{f}$ times from $\bm{p}\in\mathbb{R}^{n_{h}\times n_{w}\times d}$ to $\bm{p}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}\times d}$. Furthermore, the pre-trained DiT includes a label embedding layer, and the number of categories is 1000. Nevertheless, the used video dataset either lacks label information or encompasses a significantly smaller number of categories in comparison to ImageNet. Since we target both unconditional and class-conditional video generation, the original label embedding layer in DiT is inappropriate for our tasks, we opt to directly discard the label embedding in DiT and apply zero-initialization.
Figure 4: Temporal mask strategy.
Learning with image-video joint training. The prior work on the CNN-based video diffusion model proposes a joint image-video training strategy that greatly improves the quality of the generated videos Ho et al. (2022). We explore whether this training strategy can also improve the performance of the Transformer-based video diffusion model. To implement simultaneous training for video and image generation, We append randomly selected video frames from the same dataset to the end of the chosen videos, and each frame is independently sampled. In order to ensure our model can generate continuous videos, we propose two image-video joint training strategies: 1) tokens related to video content are used in the temporal module for modeling temporal information, while video frames tokens are excluded; 2) all tokens are used to optimize the temporal module with a new masking strategy. The first mask training strategy is that tokens related to video content are used in the temporal module to establish temporal relationships between video frames, referred to as Only Video. However, the performance is not optimal since images are only used to optimize the spatial modules. We thus propose a new training strategy that can make images also optimize the temporal modules while keeping our model to generate consistent videos, referred to as Temporal Mask. Suppose we have $f$ frames in a video clip and $i$ additional images at the end of the video, thus we have a mask matrix $M\in\mathbb{R}^{(f+i)\times(f+i)}$. The visualization of our temporal mask $M$ can be seen in Fig. 4. We evaluate the zero-shot capability on UCF101 using two different mask strategies. The FVD and FID scores for Only Video are 705.72 and 75.98, respectively, while for Temporal Mask, they are 589.86 and 63.36. These results demonstrate the superiority of Temporal Mask.
@@ 4 Experiments
4.1 Experimental setup
Datasets. We primarily conduct comprehensive experiments on four public datasets: FaceForensics Rössler et al. (2018), SkyTimelapse Xiong et al. (2018), UCF101 Soomro et al. (2012), and Taichi-HD Siarohin et al. (2019). Following the experimental setup in Skorokhodov et al. (2022), except for UCF101, we use the training split for all datasets if they are available. For UCF101, we use both training and testing splits. We extract 16-frame video clips from these datasets using a specific sampling interval, with each frame resized to 256×256 resolution for training.
Evaluation metrics. In the assessment of quantitative comparisons, we employ three evaluation metrics: Fréchet Video Distance (FVD) Unterthiner et al. (2018), Fréchet Inception Distance (FID) Parmar et al. (2021), and Inception Score (IS) Saito et al. (2017). Our primary focus rests on FVD, as its image-based counterpart FID aligns more closely with human subjective judgment. Adhering to the evaluation guidelines introduced by StyleGAN-V, we compute the FVD scores by analyzing 2,048 video clips, each comprising 16 frames. We only employ IS for assessing the generation quality on UCF101, as it leverages the UCF101-fine-tuned C3D model Saito et al. (2017).
Variant 1 Variant 2 Variant 3 Variant 4
Params (M) 673.68 673.68 676.33 676.44
FLOPs (G) 5572.69 5572.69 6153.15 1545.15
Table 1: The number of parameters and FLOPs for different model variants.
Model $N$ $D$ $H$ Param
Latte-S 12 384 6 32.48M
Latte-B 12 768 12 129.54M
Latte-L 24 1024 16 456.81M
Latte-XL 28 1152 16 673.68M
Table 2: Details of Latte models. We follow ViT and DiT configurations.
Method FaceForensics SkyTimelapse UCF101 Taichi-HD
MoCoGAN Tulyakov et al. (2018) 124.7 206.6 2886.9 -
VideoGPT Yan et al. (2021) 185.9 222.7 2880.6 -
MoCoGAN-HD Tian et al. (2021) 111.8 164.1 1729.6 128.1
DIGAN Yu et al. (2022) 62.5 83.11 1630.2 156.7
StyleGAN-V Skorokhodov et al. (2022) 47.41 79.52 1431.0 -
PVDM Yu et al. (2023) 355.92 55.41 343.6 540.2
MoStGAN-V Shen et al. (2023) 39.70 65.30 1380.3 -
LVDM He et al. (2023) - 95.20 372.0 99.0
VDT Lu et al. (2024) - - 225.7 -
W.A.L.T Gupta et al. (2024) - - 258.1 -
Latte (ours) 34.00 59.82 477.97 159.60
Latte+IMG (ours) 27.08 42.67 202.23 97.09
Table 3: FVD values of video generation models on different datasets. FVD values for other baseline models are reported and sourced from the reference StyleGAN-V or the original paper. Here, “IMG” means video-image joint training.
Methods
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
Object
Class
ModelScope Luo et al. (2023) 89.87% 95.29% 98.28% 95.79% 66.39% 52.06% 58.57% 82.25%
VideoCrafter He et al. (2023) 86.24% 92.88% 97.60% 91.79% 89.72% 44.41% 57.22% 87.34%
CogVideo Hong et al. (2022) 92.19% 95.42% 97.64% 96.47% 42.22% 38.18% 41.03% 73.40%
HiGen Qing et al. (2024) 90.07% 93.99% 93.24% 96.69% 99.17% 57.30% 63.92% 86.06%
OpenSoraPlan Lin et al. (2024) 97.27% 96.24% 99.12% 99.17% 35.28% 59.10% 65.73% 61.53%
Latte 88.88% 95.40% 98.89% 94.63% 68.89% 61.59% 61.92% 86.53%
Methods
Multiple
Object
Human
Action
Color
Spatial
Relationship
Scene
Appearance
Style
Temporal
Style
Overall
Consistency
ModelScope Luo et al. (2023) 38.98% 92.40% 81.72% 33.68% 39.26% 23.39% 25.37% 25.67%
VideoCrafter He et al. (2023) 25.93% 93.00% 78.84% 36.74% 43.36% 21.57% 25.42% 25.21%
CogVideo Hong et al. (2022) 18.11% 78.20% 79.57% 18.24% 28.24% 22.01% 7.80% 7.70%
HiGen Qing et al. (2024) 22.39% 86.20% 86.22% 22.43% 44.88% 24.54% 25.14% 27.14%
OpenSoraPlan Lin et al. (2024) 24.91% 58.20% 90.90% 49.64% 17.28% 20.04% 19.14% 20.39%
Latte 34.53% 90.00% 85.31% 41.53% 36.26% 23.74% 24.76% 27.33%
Table 4: Vbench evaluation results per dimension for different methods. A higher score demonstrates better model performance for a certain dimension. We use bold and underline to mark the best and second model performances, respectively.
Baselines. We compare with recent methods to quantitatively evaluate the outcomes, including MoCoGAN Tulyakov et al. (2018), VideoGPT Yan et al. (2021), MoCoGAN-HD Tian et al. (2021), DIGAN Yu et al. (2022), StyleGAN-V Skorokhodov et al. (2022), PVDM Yu et al. (2023), MoStGAN-V Shen et al. (2023), LVDM He et al. (2023), VDT Lu et al. (2024) and W.A.L.T Gupta et al. (2024). Furthermore, we conduct an extra comparison of IS between our proposed method and previous approaches on the UCF101 dataset.
Latte configurations. A series of $N$ Transformer blocks are used to construct our Latte model, and the hidden dimension of each Transformer block is $D$ with $N$ multi-head attention. Following ViT, we identify four configurations of Latte with different numbers of parameters as shown in Tab. 2.
Implementation details. We use the AdamW optimizer with a constant learning rate $1\times 10^{-4}$ to train all models. Horizontal flipping is the only employed data augmentation. Following common practices within generative modeling works Peebles & Xie (2023); Bao et al. (2023), an exponential moving average (EMA) of Latte weights is upheld throughout training, employing a decay rate of 0.9999. All the reported results are directly obtained from the EMA model. We borrow the pre-trained variational autoencoder from Stable Diffusion 1.4. All experiments are conducted on 8 NVIDIA A100 (80G) GPUs.
4.2 Ablation study
Video clip patch embedding. We examine the impact of two video clip patch embedding methods detailed in Sec 3.3.1. In Fig. 5a, the performance of the compression frame patch embedding method notably falls behind that of the uniform frame patch embedding method. This finding contradicts the results obtained by the video understanding method ViViT Arnab et al. (2021). We speculate that using the compression frame patch embedding method results in the loss of spatio-temporal signal, which makes it difficult for the Transformer backbone to learn the distribution of videos.
(a) Video clip patch embedding
(b) Timestep-class information injection
(c) Temporal positional embedding
(d) ImageNet pretraining
(e) Video sampling interval
(f) Model variants
Timestep-class information injection. As depicted in Fig. 5b, the performance of S-AdaLN is significantly better than that of all tokens. We believe this discrepancy may stem from the fact that all tokens only introduces timesteps or label information to the input layer of the model, which could face challenges in propagating effectively throughout the model. In contrast, S-AdaLN encodes timestep or label information into the model in a more adaptive manner for each Transformer block. This information transmission approach appears more efficient, likely contributing to superior performance and faster model convergence.
Temporal positional embedding. Fig. 5c illustrates the impact of two different temporal position embedding methods on the performance of the model. Employing the absolute position embedding approach tends to yield slightly better results than the alternative method. Since our model does not require dynamic temporal-spatial adaptability for generating variable-duration and multi-resolution videos (i.e., a key strength of RoPE), we ultimately choose absolute positional embedding as the temporal positional embedding method.
Enhancing video generation with learning strategies. As illustrated in Fig. 5d, we observe that the initial stages of training benefit greatly from the model pre-training on ImageNet, enabling rapid achievement of high-quality performance on the video dataset. However, as the number of iterations increases, the performance of the model initialized with a pre-trained model tends to stabilize around a certain level, which is far worse than that of the model initialized with random.
This phenomenon can be explained by two factors: 1) the pre-trained model on ImageNet provides a good representation, which may help the model converge quickly at an early stage; 2) there is a significant difference in data distribution between ImageNet and FaceForensics, which makes it difficult for the model to adapt the knowledge learned on ImageNet to FaceForensics. It is important to emphasize that the second point is particularly relevant for small-scale datasets (e.g., UCF101), where the domain discrepancy may outweigh the benefits of knowledge transfer. In contrast, large-scale datasets inherently exhibit greater diversity, which helps mitigate domain gaps compared to specialized small datasets. As a result, the first factor may play a more significant role in general tasks, such as text-to-video generation.
As demonstrated in Tab. 3 and Tab. 5, we find that image-video joint training (“Latte+IMG") leads to a significant improvement of FID and FVD. Concatenating additional randomly sampled frames with videos along the temporal axis enables the model to accommodate more examples within each batch, which can increase the diversity of trained models. Video sampling interval. We explore various sampling rates to construct a 16-frame clip from each training video. As illustrated in Fig. 5e, during training, there is a significant performance gap among models using different sampling rates in the early stages. However, as the number of training iterations increases, the performance gradually becomes consistent, which indicates that different sampling rates have little effect on model performance. We choose a video sampling interval of 3 to ensure a reasonable level of continuity in the generated videos to conduct the experiments of comparison to state-of-the-art.
(a) Block 0
(b) Block 7
(c) Block 13
(d) Cosine Similarity Distribution
Figure 6: Frame cosine similarity and its distribution.
Model variants. We evaluate the model variants of Latte as detailed in Sec. 3.2. We strive to equate the parameter counts across all different models to ensure a fair comparison. We commence training all the models from scratch. As shown in Fig. 5f, Variant 1 performs the best with increasing iterations. Notably, Variant 4 exhibits roughly a quarter of the floating-point operations (FLOPs) compared to the other three model variants, as detailed in Tab. 2. Therefore, it is unsurprising that Variant 4 performs the least favorably among the four variants.
Model size. We train four Latte models of different sizes according to Tab. 2 (XL, L, B, and S) on the FaceForensics dataset. Fig. 7 clearly illustrates the progression of corresponding FVDs as the number of training iterations increases. It can be clearly observed that increasing the model size generally correlates with a notable performance improvement, which has also been pointed out in image generation work Peebles & Xie (2023).
Design insights for four model variants. The design of the four variants stems from a key question: to what extent should the temporal and spatial modules be decoupled in a video generation model? While similar model structures have been employed in some previous works Gupta et al. (2024); Lu et al. (2024), they have not analyzed how these different model structures affect the video generation process. We have supplemented this with detailed experiments and analyses to help understand the impact of the degree of decoupling between the temporal and spatial modules and to explain why Variant 1 achieves the best performance.
Figure 8: The distribution of frame cosine similarity.
Excessive consecutive spatial attention modules can impair temporal coherence. As shown in Fig. 6, we conduct a detailed analysis of Variant 2, calculating the cosine similarity matrix between frame features at each block. The results demonstrate that adding spatial attention modules leads to a continuous decrease in the mean of inter-frame cosine similarity and an increase in variance. This uneven decrease in inter-frame cosine similarity ultimately results in numerous inverted relationships in the cosine similarity matrix, where, for instance, the cosine similarity between the first and second frames is lower than that between the first and fifth frames. These inverted relationships are strong evidence of the disruption of temporal coherence. We believe the root cause of this phenomenon lies in the coupled temporal-spatial information failing to achieve effective decoupling through the separation of spatiotemporal modules.
The features of the output frames from spatial attention do not match those from temporal attention. The above analysis disproves the design of stacking massive spatial attention modules, leading us to explore increasing the coupling degree between spatial and temporal modules. To this end, we design Variant 4, which modifies the alternating structure of spatiotemporal attention to a parallel structure within the Transformer block. However, experiments show that this model structure performs poorly. Analysis reveals that the performance degradation is due to the direct fusion of mismatched frame features. Specifically, as shown in Fig. 8, we visualize the distributions of inter-frame cosine similarity for the outputs of temporal attention, spatial attention, and their mutual inter-frame cosine similarity. The results show significant differences in the distributions of inter-frame cosine similarity for each type, with extremely low cosine similarity between the two. Clearly, directly adding these two types of features severely disrupts temporal information.
The design of Variant 3 is derived from TimesFormer Bertasius et al. (2021). Unlike Variant 1, which also employs an alternating spatiotemporal attention structure, Variant 3 has a higher degree of coupling. The spatial attention features are input into the temporal attention module, and only layer normalization is applied. Based on the above analysis, we believe this structure is not optimal. The simple normalization of spatial attention frame features is insufficient to make them suitable for the temporal attention layer. Experiments show that while Variant 3 performs better than Variants 2 and 4, it is weaker than Variant 1. This result further confirms the validity of our analysis.
4.3 Summary of best practices
Based on the ablation studies in Section 4.2, we identify the best practices for Transformer-based latent video diffusion models. These include model variant 1, uniform frame patch embedding, S-AdaLN, absolute position embedding, and image-video joint training. In the next section, we compare our proposed Latte, incorporating these best practices, with the current state-of-the-art.
@@ 4.3 Summary of best practices
Table 5: Inception Score and FID comparisons of Latte against other state-of-the-art on the UCF101 and FaceForensics datasets, respectively.
4.4 Comparison to state-of-the-art
Qualitative results. Fig. 9 of Appendix illustrates the video synthesis results from Latte on UCF101, Taichi-HD, FaceForensics, and SkyTimelapse. Our method consistently delivers realistic, high-resolution video generation results (256x256 pixels) in all scenarios. This encompasses capturing the motion of human faces and handling the significant transitions of athletes. Notably, our approach excels at synthesizing high-quality videos within the challenging UCF101 dataset, a task where other comparative methods often falter. More visual results are available on the project page.
Quantitative results. In Tab. 3, we provide the quantitative results of Latte and other comparative methods, respectively. Our method significantly outperforms the previous works on all datasets, which shows the superiority of our method on video generation. In Tab. 5, we report the FID on FaceForensics and the IS on UCF101 to evaluate video frame quality. Here, “IMG” means video-image joint training. Our method demonstrates outstanding performance with an FID value of 3.87 and an IS value of 73.31, significantly surpassing the capabilities of other methods.
4.5 Extension to text-to-video generation
To explore the potential capability of our proposed method, we extend Latte to text-to-video generation. We adopt the method shown in Fig. 2 (a) to construct our Latte T2V model. Sec. 4.2 mentions that leveraging pre-trained models can facilitate model training. Consequently, we utilize the weights of pre-trained PixArt-$\alpha$ ($512\times 512$ resolution) Chen et al. (2024a) to initialize the parameters of the spatial Transformer block in the Latte T2V model. Since the resolution of the commonly used video dataset WebVid-10M Bain et al. (2021) is lower than $512\times 512$, we train our model on a high-resolution video dataset Vimeo25M proposed in Wang et al. (2024a). We compare with the recent T2V models, including VideoFusion Luo et al. (2023), VideoLDM Blattmann et al. (2023b), VideoCrafter He et al. (2023), CigVideo Hong et al. (2022), HiGen Qing et al. (2024), and OpenSoraPlan Lin et al. (2024) using the quantitative metrics established by VBench Huang et al. (2024), as shown in Table 4. It demonstrates that Latte can generate comparable T2V results.
5 Conclusion
This work presents Latte, a simple and general video diffusion method, which employs a video Transformer as the backbone to generate videos. To improve the generated video quality, we determine the best practices of the proposed models, including clip patch embedding, model variants, timestep-class information injection, temporal positional embedding, and learning strategies. Comprehensive experiments show that Latte achieves state-of-the-art results across four standard video generation benchmarks. In addition, text-to-video results are comparable to those of current T2V approaches. We strongly believe that Latte can provide valuable insights for future research concerning the integration of transformer-based backbones into diffusion models for video generation as well as other modalities.
Appendix A Appendix
A.1 The sampled video frames
We provide the sampled video frames of different methods as shown in Fig. 9 of Appendix.
A.2 The structure of S-AdaLN
In Fig. 10 of Appendix, we show the structure of S-AdaLN.
A.3 Discussion about the difference from concurrent works
A similar idea has been explored in recent concurrent work VDT Lu et al. (2024), GenTron Chen et al. (2024b), W.A.L.T Gupta et al. (2024), Open-Sora Plan Lin et al. (2024), HunyuanVideo Kong et al. (2024), Pyramidal Flow Jin et al. (2025), and so on. VDT, GenTron, and W.A.L.T use an architecture akin to our Variant 3. VDT primarily focuses on generating various video tasks, including image-to-video generation and unconditional video generation, utilizing a mask learning strategy. GenTron and W.A.L.T mainly focus on general purposes, i.e., text-to-video generation and text-to-image generation. Open-Sora Plan and HunyuanVideo focus on large-scale, open-source video generation models or system frameworks, respectively. Pyramidal Flow adopts the flow matching strategy, whose core idea and training paradigm are fundamentally different from diffusion-based generative methods. Its primary goal is to improve generation efficiency. Cosmos focuses on building a world foundation model for physical AI systems, such as robotics and autonomous driving. Its goal is to simulate the dynamic behaviors of the real physical world, emphasizing physical consistency. The primary difference from the previous works is that we conduct a systematic analysis of different Transformer backbones and the relative best practices discussed in Sec. 3.2 and Sec. 3.3 on video generation.